# Mammo-Clustering:A Weakly Supervised Multi-view Global-Local Context Clustering Network for Detection and Classification in Mammography

### v1

@@ Abstract
Breast cancer has long posed a significant threat to women’s health, making early screening crucial for mitigating its impact. However, mammography, the preferred method for early screening, faces limitations such as the burden of double reading by radiologists, challenges in widespread adoption in remote and underdeveloped areas, and obstacles in intelligent early screening development due to data constraints. To address these challenges, we propose a weakly supervised multi-view mammography early screening model for breast cancer based on context clustering. Context clustering, a feature extraction structure that is neither CNN nor transformer, combined with multi-view learning for information complementation, presents a promising approach. The weak supervision design specifically addresses data limitations. Our model achieves state-of-the-art performance with fewer parameters on two public datasets, with an AUC of 0.828 on the Vindr-Mammo dataset and 0.805 on the CBIS-DDSM dataset. Our model shows potential in reducing the burden on doctors and increasing the feasibility of breast cancer screening for women in underdeveloped regions.
@@ I Introduction
Breast cancer remains the most prevalent malignant tumor among women worldwide[1]. In today’s global context, annually, more than several million women are diagnosed with breast cancer, which constitutes approximately 25% of all cancer cases among women [2]. Recent studies have highlighted that breast cancer has surpassed cardiovascular diseases as a leading cause of premature mortality globally[3] [4]. However, breast cancer is also one of the malignancies for which prevention and treatment strategies are clearly effective and efficacious[5].
Early detection is vital for reducing breast cancer mortality rates [6] [7]. Early detection can also adopt less aggressive treatment plans to improve breast retention rates and reduce the physical and psychological burden on patients. Early detection can also allow for less aggressive treatment options, reducing the physical and psychological burden on patients [8]. Additionally, it can help in identifying high-risk individuals who may benefit from preventive measures, ultimately contributing to improved overall public health outcomes.
Mammography, a low-dose X-ray technique[9], is crucial for early breast cancer detection, improving breast-conserving treatment rates and reducing mortality by over 20% [10]. It effectively identifies tumors too small to be felt, facilitating early intervention. Studies show significant mortality reduction in women aged 50 to 74 after 7 to 9 years, regardless of screening frequency [11]. Its non-invasive nature and detailed imaging enhance diagnosis and treatment planning.
However, mammography is not without its limitations, particularly concerning the risk of misdiagnosis [12]. Double reading has been proposed as a solution to reduce missed diagnoses, yet this approach significantly increases the workload for radiologists [13]. Given that the majority of mammography results are normal, the repetitive and resource-intensive nature of double reading poses a substantial burden on healthcare systems [14] [15]. Furthermore, the high costs and resource demands associated with traditional radiologist-led mammography screenings often restrict access to these services in less developed regions [16].In resource-limited settings, it is crucial to identify which interventions are most effective and feasible in reducing overall breast cancer mortality. Some researchers are dedicated to exploring hybrid screening methods that are more suitable for women in impoverished developing countries [17]. Additionally, related studies have proposed the concept of mobile mammography services [18].
In response to these challenges, substantial efforts have been made to incorporate computer-assisted detection systems to alleviate the burden on radiologists [19]. The integration of artificial intelligence (AI) for autonomous breast cancer prevention and early detection is becoming increasingly mainstream.
Relevant studies have initiated the development of breast cancer detection and classification models based on deep transfer learning [12].
Some studies have attempted to integrate digital mammography and digital breast tomosynthesis in order to detect suspected cancerous regions on mammography images [20].
However, some researchers recently collected mammographic images obtained from different breast X-ray imaging systems and investigated their performance in intelligent early breast cancer screening tasks. They found that the effectiveness of early breast cancer screening systems is often significantly constrained by the quality, accessibility, and standardization of mammographic imaging data[21] [22] [23]. Addressing these challenges is crucial for enhancing the efficiency and accessibility of breast cancer screening across diverse populations. In addition, the inherent complexity and high resolution of mammography images pose substantial challenges for AI algorithms. These images contain intricate anatomical details that are difficult for neural networks to comprehensively extract and integrate into a unified representation.
Traditional approaches to breast cancer detection and classification from mammography images predominantly rely on supervised learning techniques[24] [25] [26]. These methods typically employ object detection networks to extract feature information from mammography images, followed by a series of operations such as feature fusion, prediction, localization, and classification.
Due to the influence of mammography’s feature complexity, and breast cancer detection and classification are constrained by the quality of object detection networks, often resulting in issues such as missed detections and bounding box deviations. Moreover, the scarcity of annotated data poses significant challenges to traditional supervised learning methods in this field.
In response to these limitations, the GMIC (Global and Local Mammography Image Classification) [27] architecture introduces a novel approach based on weakly supervised learning for mammography detection and classification. This framework circumvents the constraints imposed by data scarcity, offering a promising alternative to traditional methodologies.
In traditional machine learning, each sample has a distinct label, but in multiple instance learning (MIL), samples are grouped into “bags” with known labels, while individual “instance” labels are unknown. A bag is labeled positive if at least one instance is positive; otherwise, it is negative [28]. MIL enhances generalization by training on diverse examples, reduces overfitting by exposing models to various scenarios, and improves prediction accuracy by aggregating information across instances. It is also effective for learning from weakly labeled data, useful when precise annotations are costly or difficult to obtain [29] [30].
In mammographic images, an image can be considered a “bag” containing multiple “instances,” which are different image regions or patches. Multi-Instance Learning (MIL) methods allow training with bag-level labels without precise annotations. Even if we do not know which specific instance is the lesion area, the entire bag is labeled as positive if at least one instance within it is positive [31] [27]. We can use this approach to avoid being constrained by the complexity of dataset construction. Our model draws on the concept of multi-instance learning, specifically focusing on selecting important patches and features.
Multi-view learning is a machine learning paradigm that utilizes multiple distinct sets of features, or “views,” to improve learning performance. Each view provides complementary information about the data, enhancing the model’s ability to understand complex patterns [32] [33]. Multi-view learning is employed in various domains such as image and video analysis, natural language processing, and bioinformatics. It is particularly useful in scenarios where data can be naturally divided into different perspectives [34] [35]. The primary advantages of multi-view learning include improved generalization performance, robustness to noise, and the ability to leverage complementary information from different views. By integrating diverse perspectives, multi-view learning can often achieve better accuracy and more insightful models compared to single-view approaches [32] [33]. The relevant study introduces the concept of multi-view learning into the classical generalized eigenvalue proximal support vector machine (GEPSVM) to demonstrate that multi-view learning can enhance model performance by coordinating the complementarity and consistency between different views [36]. The challenges of multi-view learning include effectively integrating information from different perspectives, as data from various viewpoints may require alignment or synchronization, which can be complex in practical applications, particularly when some perspectives may contain noise or misleading information [33].
Given that mammographic images are typically captured from different angles, such as craniocaudal (CC) and mediolateral oblique (MLO) views, and that the symmetry between the left and right breasts in the same view is considered a criterion for health in clinical diagnosis, a multi-view learning strategy is often employed. This approach leverages the complementary information from these different angles to enhance classification performance [37][38].
Convolutional networks, as the most common feature extraction paradigm, are renowned for their ability to capture spatial hierarchies and local patterns in data, making them highly popular in image processing tasks [39] [40]. However, recently, Vision Transformers (ViT) based on attention mechanisms have posed a challenge to convolutional networks [41]. By employing global self-attention operations to adaptively integrate information from patches, they have achieved impressive results.
Recent work addressing the shortcomings of these two feature extraction paradigms has significantly advanced the field of computer vision. Individuals often divide into two factions, each supporting either the convolutional or the Vision Transformer (ViT) paradigm, and focusing on their respective optimizations. Some studies have demonstrated that ResNet, with appropriate training schemes and minor modifications, can perform on par with ViT [42]. The relevant study employs an asymmetric encoder-decoder architecture to randomly mask patches of the input image and reconstruct the missing pixels, enabling efficient and effective training of large models [43]. As the competition between these two paradigms evolves, there is also a continuous integration of both [44]. Recent research, inspired by Vision Transformers (ViT), has achieved more powerful models by replacing a series of small kernels in convolution with several large convolutional kernels.
Nonetheless, beyond convolution and attention mechanisms, there are more possibilities in feature extraction tasks. The success of graph networks has also demonstrated additional potential [45]. Related work employs an anatomy-aware graph convolutional network (AGN) for early breast cancer screening in mammography, achieving promising results [46]. Recent related work directly considers the use of MLP layers for spatial interactions in the representation of image features[47].
Although clustering, as a traditional feature extraction paradigm, has gradually faded from prominence with the development of deep learning [48]. The relevant study segments the image into multiple regions by grouping a set of pixels with common characteristics. The sparsity and simplicity required for clustering demonstrate its satisfactory inter-pretability[49]. Additionally, it can connect representations of point clouds and images, reflecting its remarkable generalization capability .
We propose a Weakly Supervised Multi-instance Multi-view Mammography Classification Network, which effectively utilizes unlabeled or partially labeled data for breast cancer detection and classification.
@@
    Our work is the first to apply a non-CNN and non-attention mechanism image feature extraction method, namely Context Clustering, to the early screening task of mammography for breast cancer.
@@
    Our work achieves state-of-the-art accuracy in the early screening task of breast cancer mammography while maintaining the lowest number of parameters.
@@
    We propose a novel fusion mechanism that integrates global information, feature-based local information, and patch-based local information, placing greater emphasis on local details compared to previous methods.
@@ II Method
Our research aims to develop a rapid and reliable breast cancer early screening system to alleviate the burden on medical professionals and provide screening opportunities for women’s health in underdeveloped regions.
In this section, we describe the weakly supervised multi-instance multi-view network architecture we propose. In the first subsection, we formulate the steps involved in early breast cancer screening through mammography. In the second subsection, we provide a detailed description of the framework we propose. In the third subsection, we provide a detailed introduction and evaluation of the dataset used. In the four subsection, we introduce the evaluation metrics used to assess the effectiveness of early breast cancer screening.
@@ II-A Overall Framework
For each image $I$ in the given view, we enhance all points into 5-dimensional information points containing color and position data to obtain the set of points $S\in\mathbb{R}^{5\times w\times h}$, where $w\times h$ is equal with the number of points.
@@ II-A1 Global Information Extraction
The set $S$ is input to the first point clustering network to obtain global information $F_{g}$ and Saliency-map $I_{map}$:
$F_{g},I_{map}=f_{\textbf{global}}(S)$ (1)
Where $f_{\textbf{global}}$ represents the point clustering network for global information extraction. $F_{g}$ is extracted global information from global network. The saliency-map extracted by the global network, denoted as $I_{map}$.
$P=f_{roi}(I_{map})$
The $P$ is a set of location information, representing the positions of n example patches selected by the ROI selection function $f_{roi}$:
@@
where $p_{n}$ is a coordinate representing a position, written as $(x_{n},y_{n})$.
With $p_{n}$, we can extract n patches $\tilde{I}$ from the original image $I_{i}$ and extract n feature-based local information $F_{fl}$ from the global feature information $F_{g}$.
$F_{fl}^{n}=F_{g}(x_{n},y_{n})$ (2)
where $F_{fl}^{n}$ is means n-th feature-based local information.
II-A2 Local Information Extraction
Each selected patch $\tilde{I}_{n}$ is treated as a new image, re-enhanced based on each point to obtain its five dimensional point set $\tilde{S}_{n}$. And processed through a second point clustering network $f_{\textbf{local}}$ to obtain patch-based local information $F_{pl}^{n}$ :
$F_{pl}^{n}=f_{\textbf{local}}(\tilde{S}_{n})$ (3)
where $f_{\textbf{local}}$ represents the point clustering network for local information extraction, similar to global clustering network $f_{\textbf{global}}$.
@@ II-A3 Information Fusion and Attention Mechanism
The local information $F_{l}$ from all patches is fused with all feature-based local information and all patch-based local information :
$F_{l}=F_{fl}\oplus F_{pl}$ (4)
where Operation $\oplus$ overlays two features.
after that processed through an attention mechanism to enhance relevant features:
$F_{a}=f_{attention}(F_{l})$ (5)
Then, the attention-enhanced information is fused with the original global information, resulting in multi-instance fusion information from single-view:
$F_{f}=f_{\textbf{fuse}}(F_{g},F_{a})$ (6)
@@ II-A4 View Fusion and Classification
$F_{fusion}={f}^{\prime}_{fuse}(F_{f}^{\text{lcc}},F_{f}^{\text{lmlo}},F_{f}^{\text{rcc}},F_{f}^{\text{rmlo}})$ (7)
where $F_{f}^{\text{lcc}}$ represents the fusion feature of LCC images, other similar situations. ${f}^{\prime}_{fuse}$ is another fusion structure different from $f_{\textbf{fuse}}$. In addition to integrating features $F_{fusion}$, ${f}^{\prime}_{fuse}$ also merges $F_{g}$ and $F_{l}$ from different views.
$F_{global}={f}^{\prime}_{fuse}(F_{g}^{\text{lcc}},F_{g}^{\text{lmlo}},F_{g}^{\text{rcc}},F_{g}^{\text{rmlo}})$
$F_{local}={f}^{\prime}_{fuse}(F_{l}^{\text{lcc}},F_{l}^{\text{lmlo}},F_{l}^{\text{rcc}},F_{l}^{\text{rmlo}})$
Finally, $F_{fusion}$ performs the final classification through $f_{cls}$.
$\text{res}=f_{cls}(F_{fusion})$ (8)
@@ II-B1 From Image to Set of Points
The scale of an image can be expressed as (3, h, w), where 3 represents the RGB channels, and h and w are its height and width. We enhance each pixel by considering it as a 5-dimensional data point containing color and positional information (r, g, b, x, y). After this enhancement, the image can be represented as a set of $h\times w$ 5-dimensional data points, with a scale of $(h\times w,5)$. We can then perform feature extraction through simple clustering. From a global perspective, the image is viewed as a collection of unordered discrete data points with color and positional information. Through clustering, all points are grouped into clusters, each containing a centroid. Since each point in the set includes color and positional information, this clustering implicitly incorporates spatial and image information.
II-B2 Context Cluster Block
We employ context cluster blocks for hierarchical feature extraction, a paradigm similar to convolutional networks. At the beginning of each stage, we utilize a point reducer to decrease the number of points, thereby enhancing computational efficiency. Subsequently, a series of context cluster blocks are used to extract deep features and adaptively assign aggregated features to each uniformly selected anchor point within the cluster based on similarity, and connect and fuse the nearly points through linear projection. Finally, we perform a point-wise averaging operation on the output of the last layer.
Figure 2: The figure illustrates the clustering visualization following five steps: selecting central anchor points, identifying neighbors for each anchor, calculating features for each anchor, performing similarity analysis based on these anchors, and representing all clusters on the chart.

### v2

# Mammo-Clustering: A Weakly Supervised Multi-view Tri-level Information Fusion Context Clustering Network for Localization and Classification in Mammography

@@ Abstract
Breast cancer is a significant global health issue, and the diagnosis of breast imaging has always been challenging. Mammography images typically have extremely high resolution, with lesions occupying only a very small area. Down-sampling in neural networks can easily lead to the loss of microcalcifications or subtle structures, making it difficult for traditional neural network architectures to address these issues. To tackle these challenges, we propose a Context Clustering Network with triple information fusion. Firstly, compared to CNNs or transformers, we find that Context clustering methods (1) are more computationally efficient and (2) can more easily associate structural or pathological features, making them suitable for the clinical tasks of mammography. Secondly, we propose a triple information fusion mechanism that integrates global information, feature-based local information, and patch-based local information. The proposed approach is rigorously evaluated on two public datasets, Vindr-Mammo and CBIS-DDSM, using five independent splits to ensure statistical robustness. Our method achieves an AUC of $0.828\pm 0.020$ on Vindr-Mammo and $0.805\pm 0.020$ on CBIS-DDSM, outperforming the next best method by 3.1% and 2.4%, respectively. These improvements are statistically significant ($p<0.05$), underscoring the benefits of Context Clustering Network with triple information fusion. Overall, our Context Clustering framework demonstrates strong potential as a scalable and cost-effective solution for large-scale mammography screening, enabling more efficient and accurate breast cancer detection. Access to our method is available at https://github.com/Sohyu1/Mammo_Clustering.
Index Terms:
@@ I Introduction
Recent studies highlight that breast cancer, as the most common malignancy among women, has surpassed cardiovascular diseases to become the leading cause of premature mortality among women worldwide [1] [2]. However, breast cancer is also notably amenable to effective prevention and treatment strategies [3]. Early detection of breast cancer is crucial for reducing breast cancer mortality and significantly improving patient prognosis [4] [5]. It will allows for less invasive and more targeted treatment options, thereby reducing the physical and psychological burden on patients [6].
Mammography is a low-dose, non-invasive X-ray imaging technique [7] that plays a crucial role in the early detection of breast cancer by identifying tumors too small to be palpated, thereby facilitating timely intervention. Several studies have suggested that using mammography for early breast cancer screening can significantly reduce mortality by up to 20% [8].
One aspect of the specificity of the mammography issue is that, as a multi-view imaging technique, mammograms are typically acquired from the craniocaudal (CC) and mediolateral oblique (MLO) angles of both the left and right breasts. From a given perspective, the symmetry between the left and right breasts also serves as a critical diagnostic criterion in clinical practice. Consequently, employing a multi-view learning strategy can leverage the complementary information provided by these different imaging angles, thereby enhancing classification performance [9] [10]. Multi-view learning is a machine learning paradigm that leverages multiple feature sets, or “views,” to improve performance by capturing complementary insights. Widely applied in fields like image analysis, NLP, and bioinformatics, it enhances generalization, robustness to noise, and accuracy by integrating diverse perspectives, often outperforming single-view approaches [11][12] [13][14]. Currently, multi-view learning has become a consensus in the field of mammography analysis.
In addition, mammographic images typically possess extremely high resolution, with lesions occupying only a very small area. And handling extremely high resolution images is a challenge for traditional networks. Such as, in Convolutional Neural Networks (CNNs), layer-by-layer downsampling (like pooling) compresses feature maps, potentially leading to the loss of microcalcifications or subtle structures. Accurate localization of these lesions is crucial for diagnosis. In the case of Transformers, lesions usually occupy a very small area, and the global attention mechanism of transformers may excessively focus on irrelevant background, introducing noise. To address these limitations, several methods have been proposed to integrate global and local information to mitigate the loss of local information. The GMIC (Globally-Aware Multiple Instance Classifier) introduces an innovative framework that integrates global and local information extracted by ResNet. This concept is akin to multiple instance learning (MIL), where the selection of local information is determined by CAM to identify and extract several local regions of the image [15]. This method dynamically identifies more important regions for patch-based information extraction using a Saliency Map, then segments the original image to obtain patches. It is a weakly supervised approach that requires only image-level labels for patch-level lesion localization. This weakly supervised approach effectively reduces the model’s reliance on high-quality annotated data.
The relevant approach introduces multi-view learning based on GMIC, achieving performance improvement by integrating multi-view feature information through pooling [16].
These methods are still based on Transformer and CNN, thus the limitations of the traditional paradigms in handling high-resolution images remain unaddressed. Therefore, we opted for the COC network to process mammography images [17]. When using COC (without prior medical image training) for inference, we observed that the network inherently demonstrated a distinct intrinsic clustering capability for breast tissue and lesions, which is visually evident in Figure 1. The clustering results may be more easily correlated with anatomical or pathological features and align with clinical needs.
Figure 1: Context Clustering Visualization Diagram. Figures a to d show that the left half of each image shows the original mammogram with annotated suspicious lesions, while the right half presents Contextual clustering visualization, akin to a CNN heatmap and a VIT attention map, with the suspicious lesion locations also outlined. This figure clearly shows that the Context clustering approach effectively identifies and groups suspicious lesion areas in mammography.
Furthermore, these methods of feature fusion across different scales have not effectively utilized the information at each scale. For instance, traditional MV methods only employ global information. Conventional MIL methods, typically applied to pathological images, can only extract local information, and in GMIC framework methods, the global information extraction network extracts feature information from the entire image, but the feature information that play a role are actually few (due to the characteristic that mammography lesion areas are sparse), resulting in significant loss in the final pooled feature information, and the feature information from the global network are often wasted. And, in the current model, the local information is patch-based, lacking connections with other parts. It is merely isolated, patch-based information, which inevitably leads to the loss of global characteristics.
TABLE I: Comparison of Model Structures. There A is Multi-view structure, B is with global information, C is with patch-based local information, D is with feature-based local information.
Model A B C D
DsMIL[18]
AbMIL[19]
TransMIL[20]
SV Res[9]
SV SwinT[21]
GMIC[15]
MV Res[9]
MaMVT[10]
MV GMIC[16]
Mammo-Clustering(ours)
To address the limitations of such feature loss, we propose a three-level information fusion architecture, dividing the features when processing mammography images into global information, feature-based local information, and patch-based local information. We believe these three levels of information can maximize the utilization of mammography data. Therefore, we propose a 3-level global-local information fusion mechanism, which can be called the Tri-level Information Fusion Framework (TIFF). We will compared to previous methods (MV, MIL, GMIC), the differences are shown in the table I.
We focus on the limitation that the resolution of mammography is too high and where lesions occupy only a small portion of the image, and propose Weakly Supervised Multi-view Tri-level Information Fusion Context Clustering Network. Our primary contributions include:
@@
    This work introduces a novel non-CNN, non-attention-based feature extraction method using Context clustering for early breast cancer screening in mammography.
@@
    We propose a fusion mechanism named Tri-level Information Fusion Framework (TIFF) that integrates global, feature-based local, and patch-based local information, with enhanced focus on local details.
@@
    Our method achieves state-of-the-art accuracy with the lowest parameter count among comparable techniques, ensuring efficiency.
@@ II Method
Our research aims to design a rapid and reliable system for early breast cancer screening, with the dual objectives of reducing the workload on medical professionals and expanding access to screening opportunities for women in underdeveloped regions.
This section details the architecture of our proposed weakly supervised multi-view network. In Subsection 1, we summarize the overall workflow of our method for early breast cancer screening in mammography. Subsection 2 provides a comprehensive explanation of the key components of our framework, highlighting their individual contributions to the system’s efficacy. Finally, in Subsection 3, we introduce the loss functions utilized during training, elaborating on their roles in optimizing the network’s performance and ensuring robust learning.
@@ II-A Overall Framework
For each image $I$ in the given view, we enhance all points into 5-dimensional information points containing color and position data to obtain the set of points $S\in\mathbb{R}^{5,w\times h}$, where $w\times h$ is equal with the number of points.
@@ II-A1 Global Information Extraction
The set $S$ is input to the first Context-Clustering network to obtain global information $F_{g}$:
$F_{g}=f_{global}(S)$ (1)
where $f_{global}$ represents the Context-Clustering network for global information extraction.
The Saliency Map is obtained by processing the feature information $F_{g}$ extracted by global network through a feedforward network, denoted as $I_{map}$.
And based on the Saliency map, the ROI selection module outputs a set of location information $P$. This selection is independent of dataset annotations, and the positions of the selected n patches can be expressed as:
@@
where $p_{n}$ is a coordinate representing a position, written as $(x_{n},y_{n})$. The value of n is manually set.
With $P$, we can extract n patches $\tilde{I}$ from the original image $I$ and extract n feature-based local information $F^{n}_{fl}$ from the global information $F_{g}$. For the specified number of patches n, we recommend setting it to 4. This is because when the number of suspicious lesion locations exceeds this value, some lesions may be missed, thereby affecting the screening performance of the model. However, when the number of patches is greater than the number of suspicious lesion locations, the attention mechanism dynamically adjusts the weights of these patches, allowing the model to focus more on patches that truly contain lesion locations. Therefore, in this case, the size does not impact the model’s performance.
II-A2 Patch-based Local Information Extraction
Each selected patch $\tilde{I}_{i}$ is treated as a new image, re-enhanced based on each point to obtain its five dimensional point set $\tilde{S}_{i}$. And processed through a second Context-Clustering network to obtain patch-based local information $F_{pl}^{i}$, where $i\in\{1,2,...,n\}$.
@@ II-A3 Information Fusion and Attention Mechanism
The local information $F_{l}$ from all patches is generated by fusing all feature-based local information and all patch-based local information:
$F_{l}=f_{atten}(F_{fl}\oplus F_{pl})=f_{atten}(\sum_{i=0}^{n}F_{g}(x_{i},y_{i})\oplus\sum_{i=0}^{n}f_{local}(\tilde{S}_{i}))$ (2)
The operation $\oplus$ represents the fusion operate of the two types of information, after that processed through an attention mechanism to enhance relevant features:
Then, the local information $F_{l}$ is fused with the original global information $F_{g}$, resulting in multi-instance fusion information $F_{f}$ from single-view.
@@ II-A4 View Fusion and Classification
$F_{fusion}={f}_{fuse}(F_{f}^{\text{lcc}},F_{f}^{\text{lmlo}},F_{f}^{\text{rcc}},F_{f}^{\text{rmlo}})$ (3)
where $F_{f}^{\text{lcc}}$ represents the fusion feature of LCC images, other similar situations. ${f}_{fuse}$ is a fusion structure. In addition to integrating features $F_{fusion}$, ${f}_{fuse}$ also merges $F_{g}$ and $F_{l}$ from different views.
Finally, $F_{fusion}$ performs the final classification.
@@ II-B1 From Image to Set of Points
The scale of an image can be expressed as (3, h, w), where 3 represents the RGB channels, and h and w are its height and width. We enhance each pixel by considering it as a 5-dimensional data point containing color and positional information (r, g, b, x, y). After this enhancement, the image can be represented as a set of $h\times w$ 5-dimensional data points, with a scale of $(h\times w,5)$ [17] . We can then perform feature extraction through simple clustering. From a global perspective, the image is viewed as a collection of unordered discrete data points with color and positional information. Through clustering, all points are grouped into clusters, each containing a centroid. Since each point in the set includes color and positional information, this clustering implicitly incorporates spatial and image information.
II-B2 Context Clustering
Context Clustering performs deep feature extraction through a hierarchical structure of Context Clustering Blocks, similar to convolutional networks [17]. A point reducer is placed before each block to reduce the number of points, thereby enhancing computational efficiency. Subsequently, aggregated features are adaptively assigned to each uniformly selected anchor point within clusters based on similarity, and neighboring points are connected and fused through linear projection.
Figure 3: A visual explanation of Context-Clustering. This clustering consists of five components: selecting central anchor points, identifying neighbors for each anchor, calculating features for each anchor, performing similarity analysis based on these anchors, and representing all clusters on the chart.